OpenStudy (anonymous):

9) A rectangle has a length that is 5 feet more than its width. The area of this rectangle is 84 square feet. What is the width and length of this rectangle?

OpenStudy (anonymous):

let width be x so length will be x + 5 area = x(x+5) = 84 x^2 + 5x = 84 solve for x